Consider an acceptance sampling plan with n = 20 and c = 0. Compute the producer's risk for a lot with a defective rate of 2%.

0.6676

0.3324

0.3965

0.6035
Answer #1

Given that, an acceptance sampling plan with n = 20 and c = 0.

Defective rate = q = 2% = 0.02

=> Successive rate = p = 100 - 2 = 98% = 0.098

First we find, P(accept lot)

P(accept lot) = 20C0 * (0.98)20 * (0.02)0 = 1 * (0.98)20 * 1

=> P(accept lot) = 0.6676

Therefore,

P(producer's risk) = 1 - P(accept lot) = 1 - 0.6676 = 0.3324

=> P(producer's risk) = 0.3324